OpenStudy (chris215):

Find the specific solution of the differential equation dy/dx = 2y/x^2 with condition y(-2) = e.

OpenStudy (chris215):

I got y=e^(-2/x)

OpenStudy (freckles):

could you find the derivative to check

OpenStudy (chris215):

2e^(-2/x)/x^2?

OpenStudy (freckles):

right and e^(-2/x) is y

OpenStudy (freckles):

which gives us 2y/x^2 for y' :)

OpenStudy (freckles):

you did it @chris215

OpenStudy (freckles):

and I already check your initial condition you found the right constant

OpenStudy (chris215):

thank you:)))

OpenStudy (freckles):

np

OpenStudy (anonymous):

\[\frac{ dy }{ y }=\frac{ 2 dx }{ x^2 } \] integrating \[\ln \left| y \right|=\frac{ -2 }{ x }+c \] \[y=e ^{\frac{- 2 }{ x }+c}=e^ce ^{\frac{ -2 }{ x }}=ke ^{\frac{- 2 }{ x }} \] at x=-2 \[y=ke ^{\frac{ -2 }{ -2 }}\] \[e=ke ^{1}\] \[k=1\] \[\left| y \right|=e ^{\frac{ -2 }{ x }}\]